OpenStudy (anonymous):

HELP PLEASE!!!!! :O ---Which expression describes the rule of this sequence? (Question) 7, 14, 21, 28, 35, ? A 7+n B 7^n (Carrying the n) C 7n D n+ 7+7

OpenStudy (anonymous):

THis is an arithmetic sequence. \[a_{n} = a_{1} + (n - 1)d\] Now, plug in 7 for a_1 and 7 for d \[a_{n} = 7 + (n - 1)7\]\[a_{n} = 7 + 7n - 7\]\[a_{n} = 7n\]\THere you go :)

OpenStudy (anonymous):

np :)